1. The sum of three consecutive even numbers is 144. Find the numbers.
2. The sum of 4 consecutive odd integers is -88. Find the numbers.
3. Find three consecutive numbers such that 12 than less three times the first number is the same ad the sum of the other 2.
4.On a map 2 inches represents 35 miles. if two cities are 11 inches apart on a map,how many miles are they apart?

Answer by checkley71(8403)

X+X+2+X+4=144 OR 3X+6=144 OR 3X=138 OR X=138/3 OR X=46 X+2=48 X+4=50
PROOF 46+48+50=144 OR 144=144
X+X+2+X+4+X+6=-88 OR 4X+12=-88 OR 4X=-100 OR X=-25 X+2=-23 X+4=-21 X+6=-19
PROOF -25-23-21-19=-88 OR -88=-88
NEXT PROBLEM I CAN'T UNDERSTAND?????
2IN/35MILES=11IN/X OR 2X=35*11 OR 2X=385 OR X=385/2 OR X=192.5 MILES